# Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles with Reliable Explanations

### v4

@@ Abstract
[…]
Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles
@@
1 Introduction
A considerable proportion of published scientific articles has been found to involve fraudulent practices. It is estimated that between 1.5 and 2% of all scientific papers published in 2022 closely resemble content produced by paper-mills Noorden (2023), and about 8,000 papers were retracted from Hindawi journals due to their fraudulent origins in paper mills in 2023 Parker et al. (2024). Even more concerning, generative AI tools including those built on large language models (LLMs) can produce highly convincing fraudulent articles that have the potential to bypass existing detection mechanisms Májovskỳ et al. (2023); Kendall and Teixeira da Silva (2024). This indicates that fraud in this area is constantly adapting and learning to circumvent existing detectors Májovskỳ et al. (2023); Perkins (2023).
Fraudulence of biomedical publications is more pronounced than in other fields Grieneisen and Zhang (2012); Bik et al. (2016), with a recent finding revealing that over a fifth of newly published medical articles are at least partially fraudulent Sabel et al. (2023). Moreover, the problem of fraudulent article detection is not confined to a specific set of offending author institutions. In fact, recent high-profile cases have uncovered academic fraud from credible sources such as Harvard Medical School Johnson (2024). This widespread trend poses a serious threat to public health, as misleading medical research can directly influence clinical decisions, leading to ineffective or even harmful treatments for patients.
Prior attempts addressing this challenge mainly rely on relatively primitive heuristic techniques Parker et al. (2022); Shepperd and Yousefi (2023); Feng et al. (2024), including some based on argument quality Freedman and Toni (2024). While the threat of fraudulent articles is likely to grow in the future, the NLP community has yet to give this issue the attention it deserves. This lack of focus is partly due to two key challenges. First, there is no standard benchmark dataset for evaluating and comparing various fraud detection systems. Second, there are no open-source systems leveraging on LLMs specially designed for this task.
To address these challenges, we introduce PubMed Retraction, the first large-scale open-source benchmark for fraud detection in biomedical research, including over 11K real-world articles. Our benchmark is designed to consider the diversity across various types of academic misconduct, rather than focusing solely on specific issues like plagiarism or machine-generated text. The main goal of this benchmark is to flag potentially fraudulent articles with minimum information [^1] Complex features, such as tables and images, are not the focus of the current version of our benchmark.. Also, we release the first LLM-based system, Pub-Guard-LLM, dedicated to the task of fraud detection. Pub-Guard-LLM can be deployed in three application modes (see Figure 1): Vanilla Reasoning, Retrieval-Augmented generation, and Multi-Agent Debate, each making use of external knowledge and fine-tuning and each returning (textual) explanations of predictions. We show that Pub-Guard-LLM can significantly outperform baselines while generating more reliable explanations, by being more relevant to the explanandum Kotonya and Toni (2024) and more coherent Kotonya and Toni (2020). The three application modes provide options for users to focus on depending on their needs, balancing key factors such as precision and recall of predictions, reliability of explanations, and inference speed. This adaptability makes Pub-Guard-LLM highly versatile, accommodating a wide range of use cases in fraudulent article detection.
With the introduction of the first publicly available benchmark and an open-source LLM specifically designed for fraudulent article detection, our work aims to serve as a foundational step for future research in this field. However, while technological solutions can help address this issue, the root of the problem lies within academic evaluation systems themselves. Researchers face immense pressure to publish to secure funding and career advancements. As long as these systems prioritize metrics and publication counts over genuine scientific contributions, the arms race between fraud makers and detectors will persist.
@@ 2 Related Work
The increasing scientific fraud presents a serious challenge to both the research community and society at large. Previous studies addressing this challenge primarily focus on rule-based heuristics, based on analyzing indicators such as author affiliations Sabel et al. (2023), citation patterns Shepperd and Yousefi (2023); Feng et al. (2024), journal impact factors and author networks Pérez-Neri et al. (2022). Recently, traditional machine learning techniques Dadkhah et al. (2023); Cabanac et al. (2022) as well as BERT-based models Freedman and Toni (2024) have been proposed to help with fraudulent article detection.
However, the rapid growth in the number of fraudulent articles has not been matched by dedicated attention within the NLP field  Byrne et al. (2024). Currently, there is an urgent need for more research into two key aspects. (1) Benchmarks. Existing studies generally employ Retraction Watch The Center for Scientific Integrity (2018), a blog that monitors and reports on retractions of scientific papers, to analyze retraction behaviors Shepperd and Yousefi (2023); Byrne et al. (2022); Freedman and Toni (2024). The Retraction Watch provides a collection of retracted records, but it does not constitute a benchmark on its own for comparing existing methods. While benchmarks exist for specific fraud types, such as plagiarism Wahle et al. (2021) and machine-generated text Mosca et al. (2023); Abdalla et al. (2023); Liyanage et al. (2022), there is currently no standardized benchmark including diverse fraud types using real-world articles. (2) Open-Source Tools. While some systems for detecting fraudulent articles exist, they are often proprietary and controlled by commercial entities that are reluctant to share their technologies Christopher (2021). Even though some LLM-based applications have been developed to address news misinformation Cao et al. (2024) and plagiarism Wahle et al. (2022), there is a need for open-source LLMs specifically tailored to scientific fraud detection. Such models would enhance transparency and foster collaboration in the fight against scientific fraud.
Note that the development of open-source tools alone is not sufficient; effective fraud detection systems must also be transparent and interpretable. The necessity for systems to provide explanations for their outputs in tasks such as fact-checking Kotonya and Toni (2020) and misinformation detection Schmitt et al. (2024) is well-acknowledged. While previous research has explored incorporating explainable AI to explain decision-making processes Dalal et al. (2024); Kotonya and Toni (2024), the study of explainability in fraud detection systems remains limited.
@@
Table 1: Basic statistics of Pubmed Retraction. High Profile Rate refers to the articles deemed fraudulent despite their meta-data being ranked high (details in Appendix A).
@@ 3 PubMed Retraction Benchmark
To address the lack of a standardized benchmark for evaluating and comparing fraudulent article detection methods, we curate a large open-source benchmark for biomedical articles, PubMed Retraction, comprising over 11K real-world biomedical articles, including both legitimate and retracted publications. We select the biomedical domain due to the higher prevalence of fraudulent publications therein than in other fields  Grieneisen and Zhang (2012); Bik et al. (2016), as well as the significant impact of this type of fraud in patient safety, clinical decision-making, and public trust in medical research.
@@
To construct PubMed Retraction, we use the PubMed database Wheeler et al. (2007) and retrieve articles using the Unix command-line tool EDirect Kans (2024). To ensure relevance and avoid outdated data, we focus on articles published between 2000 and 2025. Our goal is to promote the use of minimal textual information for fraud detection, so we design the benchmark to include only title, abstract, and metadata while excluding more complex features such as full text and images. Specifically, each article in PubMed Retraction is structured to include seven key attributes: Pubmed ID, Title, Abstract, Authors, Affiliations, Journal, Is Fraudulent, with the latter a binary indicator (Yes or No). An article is labeled as fraudulent if the keyword “retracted” appears in its publication type attribute. Note that while retracted articles are not always explicitly fraudulent, they often contain errors, misconduct or unreliable findings Grieneisen and Zhang (2012), indicating potential fraud that needs to be surfaced by fraudulent article detection. To maintain a realistic distribution, we include a lower number of fraudulent articles than legitimate ones, which is aligned with real-world cases. Overall, the benchmark consists of 11,192 articles.
@@
To evaluate generalization capabilities of models, we partition articles based on keywords to create two subsets as described below. This partitioning strategy is designed for a fair assessment of a model’s ability to handle out-of-distribution samples, a critical aspect for real-world applications where models encounter unseen data. (1) General Set. This set comprises articles covering various diseases and is used for model fine-tuning. It is further split into Train, Validation, and Test subsets. (2) Cancer Set. This set consists exclusively of articles related to four types of cancer: Breast, Lung, Ovarian, and Colorectal. Articles of these four categories are excluded from the General Set, which means that these cancers are unseen diseases during fine-tuning and validation. Additionally, our benchmark includes high-profile fraudulent cases, which makes this task particularly challenging (details in Appendix A).
@@ 4 Pub-Guard-LLM
Pub-Guard-LLM is an end-to-end system for fraudulent article detection. Figure 1 provides an overview: First, the input data is enriched with external knowledge. Second, explanations to support article labels (fraudulent or not) are obtained by distilling a strong LLM (Teacher Model) via zero-shot prompting. Finally, using the article attributes and external knowledge as inputs, and distilled explanations with article labels as outputs, we fine-tune an LLM to create Pub-Guard-LLM. We propose three application modes for Pub-Guard-LLM, resulting in different performances as concerns F1, recall, reliability of explanations, and inference speed.
@@ 4.1 External Knowledge Workflow
To complement the seven input attributes (see Section 3) and incorporate task-specific knowledge, we augment external knowledge into Pub-Guard-LLM, improving LLM reasoning—a widely adopted strategy in NLP Pan et al. (2024). For this, we identify three fundamental factors to assess an article’s legitimacy: (1) author credibility, (2) affiliation reputation, and (3) journal reputation, and use the relevant external knowledge to enrich the input.
@@
In order to evaluate the credibility of authors, we use the Semantic Scholar database Lo et al. (2019), which contains millions of articles with extensive metadata and citation information, as an external knowledge source. To quantify author reputation, we use the h-index. Since LLMs are not inherently sensitive to numerical values, we map h-index scores into five reputation levels (very low, low, medium, high, very high ) using a predefined piecewise function (see Appendix A), e.g., Geoffrey Hinton (author h-index: 188, very high).
@@
As an indicator of the reputation of an author’s affiliation, we use the average citation count per institution, determined with OpenAlex Priem et al. (2022) as external knowledge source. This is a fully open index of scholarly works, which provides comprehensive data on the total publications and citations of institutions. The calculated average citation count is then mapped into five reputation levels (very low, low, medium, high, very high), e.g., Harvard University (institution average citation: 64, very high).
@@
The knowledge retrieved from external resources is integrated into the benchmark $\mathcal{D}$ for reducing duplicated queries, formally resulting in $\mathcal{D}^{\prime}=\{(x_{i},k_{i},y_{i})\}_{i=1}^{N}$ where $k_{i}$ is the external information for $(x_{i},y_{i})\in\mathcal{D}$. Note that if the inquired knowledge is missing from the external knowledge sources, we use, as $k_{i}$, the keyword “null” to indicate the absence. We choose to do so rather than omitting the field entirely, as we found it does generally indicate a lower article quality and improves the downstream performance.
@@ 4.2 Explanation Distillation
The PubMed Retraction benchmark provides only a binary label (fraudulent or not) as a target. To construct a model that classifies articles accurately while offering reliable explanations, it is essential to incorporate either human-annotated or machine-generated explanations into the ground truth. Given the high cost of human annotation, we opt for the latter as a more scalable and efficient alternative.
Recent research shows that distilling reasoning capabilities from larger (language) models can significantly enhance the performance of smaller models with fewer learning examples Li et al. (2022); Hsieh et al. (2023); Shridhar et al. (2023): teacher model rationales provide richer insights into why an input is mapped to a specific output label, capturing task-relevant knowledge missing from the input. We use a teacher model to generate explanations for each article via prompting, resulting in $\mathcal{D}^{\prime\prime}=\{(x_{i},k_{i},e_{i},y_{i})\}_{i=1}^{N}$ where $e_{i}$ is the corresponding distilled explanations with regard to $(x_{i},k_{i},y_{i})\in\mathcal{D}^{\prime}$.
Note that instruction-aligned LLMs tend to classify all articles as not fraudulent, likely to avoid controversial responses. This behavior arises because many LLMs, when fine-tuned for alignment and safety, may become overly conservative and ignore contextually relevant information Röttger et al. (2024). To mitigate this issue, it is crucial to assign a clear stance to the teacher model. For example, when evaluating a retracted article, we include the prompt: “You are tasked with providing explanations and making a firm case for why it has issues and should be retracted. You should have a clear stance.”. This ensures strong, rationale-driven explanations instead of the default neutral or overly cautious responses (see Table A2 in Appendix for an example prompt).

### v5

# Pub-Guard-LLM: Detecting Retracted Biomedical Articles with Reliable Explanations

@@ Abstract
A significant and growing number of published scientific articles ends up being retracted. Many of these retracted articles continue to be cited and influence research or clinical decisions, posing serious societal threats. In this paper, we propose Pub-Guard-LLM, the first large language model-based system tailored to retraction detection for biomedical scientific articles. We provide three application modes for deploying Pub-Guard-LLM: Vanilla Reasoning, Retrieval-Augmented Generation, and Debate, by allowing for textual explanations of prediction in each mode. To assess the performance of our system, we introduce an open-source benchmark, PubMed Retraction, comprising over 11K real-world biomedical articles, including metadata and retraction labels. We show that across all modes, Pub-Guard-LLM consistently surpasses the performance of various baselines and provides more reliable explanations, namely explanations which are deemed more relevant and coherent than those generated by the baselines when evaluated by multiple assessment methods. Pub-Guard-LLM can be used to flag potential retraction before peer review. By enhancing both detection performance and explainability in scientific retraction detection, it can contribute to reducing review workloads and preventing the spread of misinformation. The code is available at https://github.com/tigerchen52/pub_guard_llm
Pub-Guard-LLM: Detecting Retracted Biomedical Articles
@@
1 Introduction
A considerable proportion of published scientific articles are retracted. It is estimated that between 1.5 and 2% of all scientific papers published in 2022 closely resemble content from paper-mill products and retracted articles (Noorden, 2023). In 2023, about 8,000 papers were retracted from Hindawi journals due to their fraudulent origins in paper mills (Parker et al., 2024). Even more concerning, generative AI tools including those built on large language models (LLMs) can produce highly convincing articles that have the potential to bypass existing detection mechanisms (Májovskỳ et al., 2023; Kendall and Teixeira da Silva, 2024), which increases the risk of future retractions. This indicates that the area of problematic publications is constantly adapting and learning to circumvent existing detectors (Májovskỳ et al., 2023; Perkins, 2023).
Retractions of biomedical publications are more pronounced than in other fields (Grieneisen and Zhang, 2012; Bik et al., 2016), with a recent finding revealing that over a fifth of newly published medical articles have issues that may lead to retraction (Sabel et al., 2023). Moreover, the problem of retracted articles is not confined to a specific set of offending author institutions. In fact, recent high-profile cases have involved retractions from reputable organizations such as Harvard Medical School (Johnson, 2024). Many retractions happen post-publication, often after flawed papers have already influenced research or clinical decisions. Even after retraction, many articles continue to be cited without acknowledgment of their retraction Vuong (2020). This widespread trend poses a serious threat to public health, as misleading medical research can directly influence clinical decisions, leading to ineffective or even harmful treatments for patients. Therefore, it is necessary to develop a detection tool for retraction risk before peer review, which aims to flag articles that contain methodological flaws, metadata inconsistencies, or other issues of scientific integrity. Identifying these potential retractions can help stop flawed manuscripts early, ease reviewer workloads, and prevent the spread of misinformation.
Prior attempts addressing this challenge mainly rely on relatively primitive heuristic techniques (Parker et al., 2022; Shepperd and Yousefi, 2023; Feng et al., 2024), including some based on argument quality (Freedman and Toni, 2024). While the threat of retracted articles is likely to grow in the future, the AI community has yet to give this issue the attention it deserves. This lack of focus is partly due to two key challenges. First, there is no standard benchmark dataset for evaluating and comparing various retraction detection systems. Second, there are no open-source systems leveraging on LLMs specially designed for this task.
To address these challenges, we introduce PubMed Retraction, the first large-scale open-source benchmark for retraction detection in biomedical research, including over 11K real-world articles. Our benchmark is designed to consider the diversity across various types of retracted articles. The main goal of this benchmark is to flag potentially retracted articles with minimum information, such as abstracts and metadata, [^1] Complex features, such as tables and images, are not the focus of the current version of our benchmark which can help editors and conference organizers to flag potential retractions. Also, we release the first LLM-based system, Pub-Guard-LLM, dedicated to the task of retraction detection. Pub-Guard-LLM can be deployed in three application modes (see Figure 1): Vanilla Reasoning, Retrieval-Augmented generation, and Debate, each making use of external knowledge and fine-tuning and each returning (textual) explanations of predictions. We show that Pub-Guard-LLM can significantly outperform baselines while generating more reliable explanations, by being more relevant to the explanandum (Kotonya and Toni, 2024) and more coherent (Kotonya and Toni, 2020). The three application modes provide options for users to focus on depending on their needs, balancing key factors such as precision and recall of predictions, reliability of explanations, and inference speed. This adaptability makes Pub-Guard-LLM highly versatile, accommodating a wide range of use cases in retracted article detection.
With the introduction of the first publicly available benchmark and an open-source LLM specifically designed for retracted article detection, our work aims to serve as a foundational step for future research in this field. However, while technological solutions like ours can help address this issue, we believe the root of the problem lies within academic evaluation systems themselves. Researchers face immense pressure to publish to secure funding and career advancements. As long as these systems prioritize metrics and publication counts over genuine scientific contributions, the arms race between retractions and detectors will persist.
@@ 2 Related Work
The increasing number of scientific retracted papers presents a serious challenge to both the research community and society at large. Previous studies address this challenge by primarily focusing on rule-based heuristics, e.g. author affiliations (Sabel et al., 2023), citation patterns (Shepperd and Yousefi, 2023; Feng et al., 2024), journal impact factors and author networks (Pérez-Neri et al., 2022). Recently, traditional machine learning (Dadkhah et al., 2023; Cabanac et al., 2022) as well as LLM-based models (Freedman and Toni, 2024; Fletcher and Stevenson, 2025) have been proposed to help with retracted article detection.
However, the rapid growth in the number of retracted articles has not been matched by dedicated attention within our field  (Byrne et al., 2024). Currently, there is an urgent need for more research into two key aspects. (1) Benchmarks. Existing studies generally employ Retraction Watch (The Center for Scientific Integrity, 2018), a blog that monitors and reports on retractions of scientific papers, to analyze retraction behaviors (Shepperd and Yousefi, 2023; Byrne et al., 2022). The Retraction Watch provides a collection of retracted records, but it does not constitute a benchmark on its own for comparing existing methods. There is currently no standardized benchmark including diverse real-world retracted articles. (2) Open-Source Tools. While some systems for detecting retracted articles exist, they are often proprietary and controlled by commercial entities that are reluctant to share their technologies (Christopher, 2021). Even though some LLM-based applications have been developed to address news misinformation (Cao et al., 2024) and plagiarism (Wahle et al., 2022), there is a need for open-source LLMs specifically tailored to scientific retraction detection. Such models would enhance transparency and foster collaboration in the fight against problematic articles.
@@
Table 1: Basic statistics of Pubmed Retraction. High Profile Rate refers to the articles deemed retracted despite their meta-data being ranked high (details in Appendix A).
@@ 3 PubMed Retraction Benchmark
In this work, we aim to detect retraction risk, which can be defined as the identification and public flagging of a scholarly paper that contains substantial methodological flaws, metadata inconsistencies, or other issues of scientific integrity, which, if substantiated, are likely to lead to retraction.
@@
To address the lack of a standardized benchmark for evaluating and comparing retracted article detection methods, we curate a large open-source benchmark for biomedical articles, PubMed Retraction, comprising over 11K real-world biomedical articles, including both non-retracted and retracted publications. We select the biomedical domain due to the higher prevalence of retracted publications therein than in other fields  (Grieneisen and Zhang, 2012; Bik et al., 2016), as well as the significant impact of this type of retraction in patient safety, clinical decision-making, and public trust in medical research.
To construct PubMed Retraction, we use the PubMed database (Wheeler et al., 2007) and retrieve articles using the Unix command-line tool EDirect (Kans, 2024). To ensure relevance and avoid outdated data, we focus on articles published between 2000 and 2025. Our goal is to promote the use of minimal textual information for retraction detection, so we design the benchmark to include only title, abstract, and metadata. Our benchmark is intended to serve as an early warning system for decision-makers. For example, it can be used before peer review when reviews and citation data are not yet available, which enables journal editors and conference organizers to receive alerts before review assignments are made. Although we fully acknowledge that including full texts, figures, and tables could enhance this task, we aim to develop a simple yet effective solution based on minimal information (title, abstract, metadata) in this work.
@@
Specifically, each article in PubMed Retraction is structured to include seven key attributes: Pubmed ID, Title, Abstract, Authors, Affiliations, Journal, Is Retracted, with the latter a binary indicator (Yes or No). An article is labeled as retracted if the keyword “retracted” appears in its publication type attribute. To maintain a realistic distribution, we include a lower number of retracted articles than legitimate ones, which is aligned with real-world cases. Overall, the benchmark consists of 11,192 articles.
@@
To evaluate generalization capabilities of models, we partition articles based on keywords to create two subsets as described below. This partitioning strategy is designed for a fair assessment of a model’s ability to handle out-of-distribution samples, a critical aspect for real-world applications where models encounter unseen data. (1) General Set. This set comprises articles covering various diseases and is used for model fine-tuning. It is further split into Train, Validation, and Test subsets. (2) Cancer Set. This set consists exclusively of articles related to four types of cancer: Breast, Lung, Ovarian, and Colorectal. Articles of these four categories are excluded from the General Set, which means that these cancers are unseen diseases during fine-tuning and validation. Additionally, our benchmark includes high-profile retracted cases, which makes this task particularly challenging (details in Appendix A).
@@ 4 Pub-Guard-LLM
Pub-Guard-LLM is an end-to-end system for retracted article detection. Figure 1 provides an overview. First, the input data is enriched with external knowledge. Second, explanations to support article labels (retracted or not) are obtained by distilling a strong LLM (Teacher Model) via zero-shot prompting. Finally, using the article attributes and external knowledge as inputs, and distilled explanations with article labels as outputs, we fine-tune an LLM to create Pub-Guard-LLM. We propose three application modes for Pub-Guard-LLM, resulting in different performances as concerns F1, recall, reliability of explanations, and inference speed.
@@ 4.1 External Knowledge Workflow
To complement the seven input attributes (see Section 3) and incorporate task-specific knowledge, we augment external knowledge into Pub-Guard-LLM, improving LLM reasoning—a widely adopted strategy in NLP (Pan et al., 2024). For this, we identify three fundamental factors to assess an article’s legitimacy: (1) author credibility, (2) affiliation reputation, and (3) journal reputation, and use the relevant external knowledge to enrich the input.
@@
In order to evaluate the credibility of authors, we use the Semantic Scholar database (Lo et al., 2019), which contains millions of articles with extensive metadata and citation information, as an external knowledge source. To quantify author reputation, we use the h-index. Since LLMs are not inherently sensitive to numerical values, we map h-index scores into five reputation levels (very low, low, medium, high, very high) using a predefined piecewise function (see Appendix A), e.g., Geoffrey Hinton (author h-index: 188, very high).
@@
As an indicator of the reputation of an author’s affiliation, we use the average citation count per institution, determined with OpenAlex (Priem et al., 2022) as external knowledge source. This is a fully open index of scholarly works, which provides comprehensive data on the total publications and citations of institutions. The calculated average citation count is then mapped into five reputation levels (very low, low, medium, high, very high), e.g., Harvard University (institution average citation: 64, very high).
@@
The knowledge retrieved from external resources is integrated into the benchmark $\mathcal{D}$ for reducing duplicated queries, formally resulting in $\mathcal{D}^{\prime}=\{(x_{i},k_{i},y_{i})\}_{i=1}^{N}$ where $k_{i}$ is the external information for $(x_{i},y_{i})\in\mathcal{D}$. Note that, if the inquired knowledge is missing from the external knowledge sources, we use, as $k_{i}$, the keyword “null” to indicate the absence. We choose to do so rather than omitting the field entirely, as we found it does generally indicate a lower article quality and improves the downstream performance.
@@ 4.2 Explanation Distillation
The PubMed Retraction benchmark provides only a binary label (retracted or not) as a target. To construct a model that classifies articles accurately while offering reliable explanations, it is essential to incorporate either human-annotated or machine-generated explanations into the ground truth. Given the high cost of human annotation, we opt for the latter as a more scalable and efficient alternative.
Recent research shows that distilling reasoning capabilities from larger (language) models can significantly enhance the performance of smaller models with fewer learning examples (Li et al., 2022; Hsieh et al., 2023; Shridhar et al., 2023): teacher model rationales provide richer insights into why an input is mapped to a specific output label, capturing task-relevant knowledge missing from the input. What is more interesting is that even if some intermediate steps of the distilled explanations contain inaccuracies, student models can still have enhancements by learning reasoning flows Li et al. (2023); Hsieh et al. (2023). We use a teacher model to generate explanations for each article via prompting, resulting in $\mathcal{D}^{\prime\prime}=\{(x_{i},k_{i},e_{i},y_{i})\}_{i=1}^{N}$ where $e_{i}$ is the corresponding distilled explanations with regard to $(x_{i},k_{i},y_{i})\in\mathcal{D}^{\prime}$.
Note that instruction-aligned LLMs tend to classify all articles as not retracted, likely to avoid controversial responses. This behavior arises because many LLMs, when fine-tuned for alignment and safety, may become overly conservative and ignore contextually relevant information (Röttger et al., 2024). To mitigate this issue, it is crucial to assign a clear stance to the teacher model. For example, when evaluating a retracted article, we include the prompt: “You are tasked with providing explanations and making a firm case for why it has issues and should be retracted. You should have a clear stance”. This ensures strong, rationale-driven explanations instead of the default neutral or overly cautious responses (see Table A3 in Appendix for an example prompt). Additionally, we conduct human evaluations to validate the coherence and relevance of distilled explanations (see Section C.2 in Appendix).